# Gait Transitions in Load-Pulling Quadrupeds: Insights from Sled Dogs and a Minimal SLIP Model

### v2

@@ Abstract
Quadrupedal animals employ diverse galloping strategies to optimize speed, stability, and energy efficiency. However, the biomechanical mechanisms that enable adaptive gait transitions during high-speed locomotion under load remain poorly understood. In this study, we present new empirical and modeling insights into the biomechanics of load-pulling quadrupeds, using sprint sled dogs as a model system. High-speed video and force recordings reveal that sled dogs often switch between rotary and transverse galloping gaits within just a few strides and without any observable changes in speed, stride duration, or terrain, providing clear evidence of locomotor multistability during high-speed load-pulling. To investigate the mechanical basis of these transitions, a physics-based quadrupedal Spring-Loaded Inverted Pendulum model with hybrid dynamics and prescribed footfall sequences to reproduce the asymmetric galloping patterns observed in racing sled dogs. Through trajectory optimization, we replicate experimentally observed gait sequences and identify swing-leg stiffness modulation as a key control mechanism for inducing transitions. This work provides a much-needed biomechanical perspective on high-speed animal draft and establishes a modeling framework for studying locomotion in pulling quadrupeds, with implications for both biological understanding and the design of adaptive legged systems.
@@ I Introduction
Coordinated limb movement is fundamental to effective animal locomotion and has long inspired the design of agile, efficient robotic systems [1, 2]. Among terrestrial mammals, quadrupedal galloping is one of the most mechanically dynamic and energetically demanding gait patterns. Species such as dogs, horses, and cheetahs adopt distinct galloping strategies—particularly rotary and transverse variants—to optimize performance across different speeds and terrains [3, 4]. Traditionally viewed as a stereotyped, high-speed gait with transitions linked to velocity or environmental changes [5], galloping has more recently been reinterpreted through the lens of locomotor multistability, where multiple gait solutions can coexist under nearly identical conditions [6]. For instance, modeling studies have shown that even small variations in limb compliance or body symmetry can give rise to distinct galloping modes within the same speed range [7, 8]. However, such phenomena have rarely been examined in quadrupeds operating under external load. Sled dogs offer a particularly compelling and under-studied example of this phenomenon: they sustain high-speed galloping while pulling a sled, requiring coordination not only for self-propulsion but also to generate forward force along a connecting rope known as the tugline. This added load introduces unique biomechanical constraints, as the resulting tension likely affects stance duration, limb loading, and possibly interlimb timing.
@@
Despite growing interest in quadrupedal locomotion, particularly in working contexts, the dynamics of load-pulling quadrupeds remain poorly understood. A long history of animal draft in agriculture, transportation, and sport has generated limited biomechanical literature on load-pulling quadrupeds, especially under dynamic conditions. Most existing work has focused on practical questions in agricultural settings, emphasizing work output or harness efficiency during heavy-load, static-pulling scenarios [9, 10, 11, 12, 13]. Recent robotics research has begun to model dynamic quadruped-load systems [14], but such models typically operate at the path-planning level and do not capture limb-level dynamics or the energetic coupling between the quadruped and its load. To date, no dynamical model has been developed that captures the limb-load coupling or stride-level coordination necessary to replicate the complex dynamics of quadruped-load systems, especially those involving gait transitions.
@@
In this study, we contribute new empirical and modeling insights to the study of load-pulling quadrupeds. Using high-speed video recordings of racing sled dogs, we reconstructed footfall sequences and tugline forces during high-speed locomotion. Our analysis reveals that sprint sled dogs do not rely on a single fixed gallop. Instead, they frequently switch between rotary and transverse galloping patterns, often within a few strides and without apparent external triggers, demonstrating a high degree of locomotor multistability.
@@
We build on the Spring-Loaded Inverted Pendulum (SLIP) framework [15, 16, 17], a reduced-order model widely used to study locomotor dynamics, and extend it to quadrupedal systems that incorporate limb asymmetry, external load pulling, and predefined footfall sequences. While prior models have explored gait energetics [18] and bipedal transitions [19], our formulation captures the full-body dynamics and interaction forces between a quadruped and a load. In contrast to neuromechanical simulations that require detailed anatomical modeling [20], our model achieves generalizability with fewer parameters while remaining grounded in physical dynamics. To our knowledge, this is the first dynamical model capable of reproducing not only the detailed stride mechanics but also the spontaneous gait transitions observed in load-bearing quadrupeds. By embedding empirical gait sequences and optimizing against experimental data, our model provides new insight into the control strategies and mechanical feasibility of stride-to-stride transitions in pulling contexts.
@@ II.1 Gait Pattern Identification
Fieldwork was conducted in collaboration with a professional sprint-racing sled dog kennel, with data collection efforts integrated into their regular training routine. Sprint-racing is a fast-paced style of dogsled racing, in which dogs typically maintain galloping speeds of 25 to 40 km per hour over distances less than 50 km [21]. Training runs took place on dirt trails or shallow snow, with teams pulling a wheeled training rig carrying one to two human passengers.
@@
Precise measurement of limb-specific stance and stride timing typically relies on controlled experiments using high-speed video, with either manual annotation [22, 23] or automated labeling methods [24]. However, replicating such methods in the field, especially with free-running sled dogs, is logistically challenging. To overcome this, we mounted an action camera (GoPro, 60 fps +) on an extended pole held from the training rig. This setup enabled us to capture targeted side-view footage of two individuals—referred to here as Individual 1 and Individual 2—during select training runs. From this footage, we manually annotated stance and stride durations for each limb (Fig. 1A).
@@
These annotations were used to classify gait patterns based on the unique sequence of limb-ground interactions. Following Hildebrand’s convention [4], a footfall diagram can be used to depict contact durations as solid bars and aerial phases as gaps. A short selection of footfalls, presented this way (Fig. 1B), illustrates the frequent switching between transverse and rotary gallops during high-speed pulling. In transverse gallop, the hind legs strike in sequence, followed by the contralateral and then ipsilateral forelegs. In rotary gallop, the placement of the second hind foot is followed by the ipsilateral forefoot, creating a seemingly rotating sequence of footfalls around the body [3]. All four permutations of these galloping patterns are illustrated in Fig. 2.
@@
Our analysis also benefits from data collected in parallel with video using custom data loggers integrated into the dogs’ harnesses. These sensors were clipped onto a standard ‘X-back’ harness near the dog’s withers, and measured tugline force (62 Hz) and tri-axial acceleration (125 Hz) throughout each stride. After experiments were completed, data was transferred from the data loggers to a computer for post-hoc analysis. For this paper, only the force data was analyzed. For a more detailed overview of the sensors used, including design files and firmware, see  [^1] https://github.com/DLARlab/2025 Gait Transitions in Load Pulling Quadrupeds Insights.git. The GoPro camera recorded GPS data in the video file metadata, and the data logger timestamped samples using an included GPS module. The shared epoch timestamps allowed force measurements to be synchronized with video recordings. Fig. 1B shows this tension over a short period of time. Notably, the measured force exhibits stride-synchronized oscillations, with peaks that often—but not always—coincide with the touch-down of the trailing forelimb. Load cells were calibrated between runs using fixed weights to correct for mechanical drift or deformation.
@@ II.2 Gait Pattern Definitions
Figure 2: Quadrupedal footfall patterns following Hildebrand’s convention [4]. The horizontal axis represents normalized stride time. Colored bars denote ground contact phases for each leg: green (LH), red (LF), yellow (RF), and blue (RH). Each row represents a distinct galloping gait, with white backgrounds for transverse sequences and grey for rotary. Touchdown ($t_{i}^{\scriptstyle\text{TD}}$) and liftoff ($t_{i}^{\scriptstyle\text{LO}}$) define the stance period. Phase lags are annotated for fore-hind ($d_{p}$), hindlimb ($d_{h}$), and forelimb ($d_{f}$) timing. Below each bar plot, cartoons depict limb configurations, and adjacent dot plots show footfall sequences, with green and blue font labels for left- and right-leading gaits.
@@
To standardize comparisons across gait types, we define an extended flight phase between the LO of the trailing hindlimb and the TD of the leading forelimb, during which the torso is airborne. The start of each stride is anchored to the midpoint of this aerial phase, corresponding to the moment of peak torso height. This convention follows biomechanical interpretations of galloping dynamics and ensures a consistent phase reference for stride analysis [7, 26]. Each stride is encoded by the timing vector:
@@
Phase lags of zero correspond to bounding or half-bounding gaits, while values of $0.5$ indicate symmetrical gaits such as pacing or trotting, following Hildebrand’s classification [4]. Within this space, each canonical galloping gait $G_{e}$, for $e\in\{\textit{TranL},\textit{TranR},\textit{RotL},\textit{RotR}\}$, is uniquely determined by the sign pair $(d_{f},d_{h})$:
@@ III.1 Quadrupedal SLIP Model with Load
We extend a planar SLIP model to capture the dynamics of sled dogs pulling a load at high speed. The base model, adapted from our previous work [27, 7], consists of a rigid torso supported by four massless spring legs and augmented with a discrete load mass connected via a unilateral spring element that mimics the intermittent tension of a tugline.
@@
The quadruped consists of a rigid body of mass $M_{\text{torso}}$ and moment of inertia $J_{\text{torso}}$, supported by four limbs anchored at fixed hip and shoulder joints. The system evolves in the sagittal plane and is described by eight generalized coordinates: the torso position $(q_{x},q_{z})$, pitch angle $q_{\theta}$, four leg swing angles $q_{\alpha,i}$ for $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$, and the relative horizontal displacement of the load $q_{\text{load}}$.
@@
The model forms a hybrid dynamical system, with stance and flight phases governed by distinct sets of equations. The hind and fore legs are mounted at distances $l_{b,H}$ and $l_{b,F}$ from the torso Center of Mass (CoM), respectively. In stance, ground contact imposes holonomic constraints and generates ground reaction forces (GRFs); in flight, limbs follow unconstrained passive dynamics driven by torsional stiffness. Each leg behaves as a linear spring with stiffness $k_{\text{leg}}$ during stance, storing and releasing energy to propel the body forward. Legs are massless but terminate in point feet with small masses $m_{\text{foot}}$ to define discrete ground contact events [28]. A fore–hind stiffness ratio $r_{\text{leg}}$ is included to model anatomical asymmetry between limb pairs. During flight, each leg swings passively under a torsional spring of stiffness $k_{\text{swing}}$, resulting in approximately sinusoidal motion.
@@
To represent the external load, we introduce a point mass $M_{\text{load}}$ connected to the torso by a unilateral, massless spring (the tugline) with stiffness $k_{\text{tug}}$ and resting length $l_{t}$. The tugline exerts force only when stretched beyond its rest length, producing discontinuous coupling between the quadruped and the load. When slack, the load follows its inertial trajectory; when taut, it exerts a horizontal restoring force that influences both the load and torso dynamics. To maintain steady galloping in the presence of energy loss (e.g., due to sliding friction at the load), similar to the Inverted Pendulum (IP) model with rigid legs [29], we tilt the terrain by a small slope angle $\theta_{\text{slope}}$ and model Coulomb friction at the load-ground interface with coefficient $\mu_{\text{ground}}$. This slope enables gravitational work to balance energy dissipation, facilitating sustained periodic motion. All parameters are summarized in Table 1 and are normalized by the total system mass $m$, gravitational acceleration $g$, and resting leg length $l_{0}$.
@@
Table 1: Key model parameters used in the quadrupedal SLIP model with load. All quantities are normalized using total mass $m$, gravity $g$, and resting leg length $l_{0}$ [28].
Symbol Description Units
$M_{\rm torso}$ Torso mass $m$
$m_{\rm foot}$ Mass of each foot $m$
$J_{\rm torso}$ Torso pitching inertia $ml_{0}^{2}$
$k_{\rm leg}$ Leg spring stiffness $mg/l_{0}$
$r_{\rm leg}$ Stiffness ratio between hind/fore legs $\cdot$
$k_{\mathrm{swing},i}^{n}$ Torsional stiffness of swing‐leg spring $mgl_{0}$
$l_{l}$ Resting leg length $l_{0}$
$l_{b,H}$, $l_{b,F}$ Distance from CoM to hip/shoulder $l_{0}$
$M_{\rm load}$ Load mass $m$
$k_{\rm tug}$ Tugline (unilateral spring) stiffness $mg/l_{0}$
$l_{t}$ Tugline resting length $l_{0}$
$\mu_{\text{ground}}$ Friction coefficient of ground contact $\cdot$
$\theta_{\text{slope}}$ Slope angle of running surface $\cdot$

### v4

@@ Abstract
Quadrupedal animals employ diverse galloping strategies to optimize speed, stability, and energy efficiency. However, the biomechanical mechanisms that enable adaptive gait transitions during high-speed locomotion under load remain poorly understood. In this study, we present new empirical and modeling insights into the biomechanics of load-pulling quadrupeds, using sprint sled dogs as a model system. High-speed video and force recordings reveal that sled dogs often switch between rotary and transverse galloping gaits within just a few strides and with minimal changes in speed and stride duration, suggestive of of locomotor multistability during high-speed load-pulling. To investigate the mechanical basis of these transitions, a physics-based quadrupedal Spring-Loaded Inverted Pendulum model with hybrid dynamics and prescribed footfall sequences to reproduce the asymmetric galloping patterns observed in racing sled dogs. Through trajectory optimization, we replicate experimentally observed gait sequences and identify swing-leg stiffness modulation as a key control mechanism for inducing transitions. This work provides a much-needed biomechanical perspective on high-speed animal draft and establishes a modeling framework for studying locomotion in pulling quadrupeds, with implications for both biological understanding and the design of adaptive legged systems.
@@ I Introduction
Coordinated limb movement is fundamental to effective animal locomotion biewener2003animal; dickinson2000animal. Among terrestrial mammals, quadrupedal galloping is one of the most mechanically dynamic and energetically demanding gait patterns. Species such as dogs, horses, and cheetahs adopt distinct galloping strategies—particularly rotary and transverse variants—to optimize performance across different speeds and terrains biancardi2012biomechanical; Hildebrand1977. Traditionally, galloping has been viewed as a stereotyped high-speed gait, with transitions primarily linked to locomotor speed and related energetics Hoyt1981gaitenergetic. More recent work in biomechanics and legged robotics suggests that gait preferences and transitions are not governed by speed alone, but can instead emerge from feedback-driven responses to stability demands, external loading, and terrain-dependent perturbations shafiee2024viability; Farley1991trigger; owaki2017quadruped; zhang2024learning; wilshin2017longitudinal. Lead limb switching within asymmetrical gaits has also been proposed as a wear-leveling strategy to redistribute limb-specific demand and mitigate fatigue walter2007ground. These perspectives motivate a reinterpretation of galloping (and other gaits) through the lens of locomotor multistability, in which multiple gait solutions can coexist under nearly identical conditions and be selectively expressed in response to perturbations or control objectives wilshin2020dog. For instance, modeling studies have shown that even small variations in limb compliance or body symmetry can give rise to distinct galloping modes within the same speed range Ding2024RALSymmetry; alqaham202516. Despite growing interest in context-dependent gait selection and switching, stride-resolved studies of high-speed galloping under sustained external load (e.g., harness-mediated pulling) remain rare. Sled dogs offer a particularly compelling and under-studied example of this phenomenon: they sustain high-speed galloping while pulling a sled, requiring coordination not only for self-propulsion but also to generate forward force along a connecting rope known as the tugline. This added load introduces unique biomechanical constraints, as the resulting tension likely affects limb loading and possibly interlimb timing.
@@
Although our study focuses on biological quadrupeds, the mechanisms uncovered here are also relevant for the design and control of agile, load-pulling quadrupedal robots. Despite growing interest in quadrupedal locomotion, particularly in working contexts, the dynamics of load-pulling quadrupeds remain poorly understood. A long history of animal draft in agriculture, transportation, and sport has generated limited biomechanical literature on load-pulling quadrupeds, especially under dynamic conditions. Most existing work has focused on practical questions in agricultural settings, emphasizing work output or harness efficiency during heavy-load, static-pulling scenarios starkey1989animal; starkey1989harnessing; lawrence1993experimental; Rooney1985; Bukhari2023. Recent robotics research has begun to model dynamic quadruped-load systems RobotLoadPathPlaning2024, but such models typically operate at the path-planning level and do not capture limb-level dynamics or the energetic coupling between the quadruped and its load. To the best of our knowledge, existing dynamical models of quadruped–load systems do not simultaneously capture limb–load coupling and stride-level coordination in a way that reproduces the experimentally observed gait transitions.
@@
In this study, we contribute new empirical and modeling insights to the study of load-pulling quadrupeds. Using high-speed video recordings of racing sled dogs, we reconstructed footfall sequences and tugline forces during high-speed locomotion. Our analysis reveals that sprint sled dogs do not rely on a single fixed gallop. Instead, they frequently switch between rotary and transverse galloping patterns, often within a few strides and without large changes in forward speed, consistent with a multistable gait repertoire.
@@
We build on the Spring-Loaded Inverted Pendulum (SLIP) framework geyer2006spring; full1999templates; fukuhara2018spontaneous, a reduced-order model widely used to study locomotor dynamics, and extend it to quadrupedal systems that incorporate limb asymmetry, external load pulling, and predefined footfall sequences. While prior models have explored gait energetics alqaham2024sixteen and bipedal transitions DingJerboa2022, our formulation captures the full-body dynamics and interaction forces between a quadruped and a load. In contrast to neuromechanical simulations that require detailed anatomical modeling geyer2010muscle, our model achieves generalizability with fewer parameters while remaining grounded in physical dynamics. To our knowledge, this is the first dynamical model capable of reproducing not only the detailed stride mechanics but also the spontaneous gait transitions observed in load-bearing quadrupeds. By embedding empirical gait sequences and optimizing against experimental data, we show that these transitions can be generated primarily through targeted modulation of limb torsional stiffness, with hind-limb stiffness acting as a dominant control-like parameter and forelimb stiffness enabling changes in leading-leg sequence. This identifies differential limb stiffness as a simple, mechanically grounded strategy for organizing stride-to-stride transitions in quadruped–load systems and delineates the region of parameter space in which such transitions are feasible.
@@ II.1 Gait Pattern Identification
Fieldwork was conducted in collaboration with a professional sprint-racing sled dog kennel, with data collection efforts integrated into their regular training routine. Sprint-racing is a fast-paced style of dogsled racing, in which dogs typically maintain galloping speeds of 25 to 40 km per hour over distances less than 50 km thorsrud2021description. Training runs took place on dirt forest roads or shallow snow, with teams pulling a motorized ATV carrying two human passengers. We restricted analysis to straight, steady-running segments. Specifically, we excluded intervals containing starts/stops, overt turns, or obvious driver-induced speed changes. Throughout, the driver aimed to maintain race-like speeds, but speed and acceleration were not strictly fixed.
@@
Precise measurement of limb-specific stance and stride timing typically relies on controlled experiments using high-speed video, with either manual annotation Hilderbrand1989Quadrupedal; sandberg2020review or automated labeling methods sheppard2022stride. However, replicating such methods in the field, especially with free-running sled dogs, is logistically challenging. To overcome this, we mounted an action camera (GoPro, 60 fps +) on an extended pole held from the training rig. This setup enabled us to capture targeted side-view footage of two individuals—referred to here as Individual 1 and Individual 2—during select training runs. From this footage, we manually annotated stance and stride durations for each limb (Fig. 1A).
@@
These annotations were used to classify gait patterns based on the unique sequence of limb-ground interactions. Following Hildebrand’s convention Hildebrand1977, a footfall diagram can be used to depict contact durations as solid bars and aerial phases as gaps. A short selection of footfalls, presented this way (Fig. 1B), illustrates the frequent switching between transverse and rotary gallops during high-speed pulling. In transverse gallop, the hind legs strike in sequence, followed by the contralateral and then ipsilateral forelegs. In rotary gallop, the placement of the second hind foot is followed by the ipsilateral forefoot, creating a seemingly rotating sequence of footfalls around the body biancardi2012biomechanical. All four permutations of these galloping patterns are illustrated in Fig. 2.
@@
Our analysis also benefits from data collected in parallel with video using custom data loggers integrated into the dogs’ harnesses. These sensors were clipped onto a standard ‘X-back’ harness near the dog’s withers, and measured tugline force (62+ Hz) and tri-axial acceleration (120+ Hz) throughout each stride. After experiments were completed, data was transferred from the data loggers to a computer for post-hoc analysis. For this paper, only the force data was analyzed. For a more detailed overview of the sensors used, including design files and firmware, see the Data and materials availability section. The GoPro camera recorded GPS data with absolute (epoch) timestamps in the video file metadata, and the data logger used an onboard GPS module to timestamp each sensor sample (force and acceleration) on the same epoch time base. During analysis, we synchronized the video frames with the logger data by aligning these GPS timestamps.Fig. 1B shows this tension over a short period of time. Notably, the measured force exhibits stride-synchronized oscillations, with peaks that often—but not always—coincide with the touch-down of the trailing forelimb. Load cells were calibrated between runs using fixed weights to correct for mechanical drift or deformation.
Dog stride frequency was computed stride by stride as the time between consecutive touchdowns of the trailing hind limb, which approximates the full galloping cycle. While the animal-borne sensors had GPS, these were used solely for temporal synchronization. Forward speed was instead estimated from a separate GPS receiver mounted on the ATV. Latitude and longitude were converted to planar coordinates, and instantaneous ground speed was computed as the magnitude of the central-difference displacement divided by the corresponding time interval and then smoothed with a 5-s centered moving-average filter. Because the dogs were continuously tethered to the ATV along straight, approximately level road segments, we treated ATV speed as the forward speed of the team. The resulting time series of stride frequency, forward speed, and timing of identified gait transitions are shown in Fig. 4.
@@ II.2 Gait Pattern Definitions
Figure 2: Quadrupedal footfall patterns following Hildebrand’s convention Hildebrand1977 showing (A) Transverse galloping, left leading (TranL), (B) Transverse galloping, right leading (TranR), (C) Rotary galloping, left leading (RotL), (D) Rotary galloping, right leading (RotR). The horizontal axis represents normalized stride time. Colored bars denote ground contact phases for each leg: green (LH), red (LF), yellow (RF), and blue (RH). Each row represents a distinct galloping gait, with white backgrounds for transverse sequences and grey for rotary. Touchdown ($t_{i}^{\scriptstyle\text{TD}}$) and liftoff ($t_{i}^{\scriptstyle\text{LO}}$) define the stance period. Phase lags are annotated for fore-hind ($d_{p}$), hindlimb ($d_{h}$), and forelimb ($d_{f}$) timing. Below each bar plot, cartoons depict limb configurations, and adjacent dot plots show footfall sequences, with green and blue font labels for left- and right-leading gaits.
@@
To standardize comparisons across gait types, we define an extended flight phase between the LO of the trailing hindlimb and the TD of the leading forelimb, during which the torso is airborne. The start of each stride is anchored to the midpoint of this aerial phase, corresponding to the moment of peak torso height. This convention follows biomechanical interpretations of galloping dynamics and ensures a consistent phase reference for stride analysis Ding2024RALSymmetry; DingJerboa2022. Each stride is encoded by the timing vector:
@@
Phase lags of zero correspond to bounding or half-bounding gaits, while values of $0.5$ indicate symmetrical gaits such as pacing or trotting, following Hildebrand’s classification Hildebrand1977. Within this space, each canonical galloping gait $G_{e}$, for $e\in\{\textit{TranL},\textit{TranR},\textit{RotL},\textit{RotR}\}$, is uniquely determined by the sign pair $(d_{f},d_{h})$:
@@ III.1 Quadrupedal SLIP Model with Load
We extend a planar SLIP model to capture the dynamics of sled dogs pulling a load at high speed. The base model, adapted from our previous work GanDynamicSimilarity; Ding2024RALSymmetry, consists of a rigid torso supported by four massless spring legs and augmented with a discrete load mass connected via a unilateral spring element that mimics the intermittent tension of a tugline.
@@
The quadruped consists of a rigid body of mass $M_{\text{torso}}$ and moment of inertia $J_{\text{torso}}$, supported by four limbs anchored at fixed hip and shoulder joints. The system evolves in the sagittal plane and is described by eight generalized coordinates: the torso position $(q_{x},q_{z})$, pitch angle $q_{\theta}$, four leg swing angles $q_{\alpha,i}$ for $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$, and the relative horizontal displacement of the load $q_{\text{load}}$. Because the model is constrained to sagittal-plane motion and does not include limb-specific costs or fatigue dynamics, it is effectively symmetric with respect to left- versus right-leading solutions.
@@
The model forms a hybrid dynamical system, with stance and flight phases governed by distinct sets of equations. The hind and fore legs are mounted at distances $l_{b,H}$ and $l_{b,F}$ from the torso Center of Mass (CoM), respectively. In stance, ground contact imposes holonomic constraints and generates ground reaction forces (GRFs); in flight, limbs follow unconstrained passive dynamics driven by torsional stiffness. Each leg behaves as a linear spring with stiffness $k_{\text{leg}}$ during stance, storing and releasing energy to propel the body forward. Legs are massless but terminate in point feet with small masses $m_{\text{foot}}$ to define discrete ground contact events GanAllCommonBipedal. A fore–hind stiffness ratio $r_{\text{leg}}$ is included to model anatomical asymmetry between limb pairs. During flight, each leg swings passively under a torsional spring of stiffness $k_{\text{swing}}$, resulting in approximately sinusoidal motion. Biologically, $k_{\text{swing}}$ represents an effective hip/shoulder swing impedance modulated by muscle activation, while $k_{\text{leg}}$ represents an effective stance-phase limb stiffness arising from musculoskeletal compliance and activation.
@@
To represent the external load, we introduce a point mass $M_{\text{load}}$ connected to the torso by a unilateral, massless spring (the tugline) with stiffness $k_{\text{tug}}$ and resting length $l_{t}$. The tugline exerts force only when stretched beyond its rest length, producing discontinuous coupling between the quadruped and the load. When slack, the load follows its inertial trajectory; when taut, it exerts a horizontal restoring force that influences both the load and torso dynamics. To maintain approximately steady galloping in the presence of energy loss (e.g., due to sliding friction at the load), similar to the inverted pendulum model with rigid legs kuo2005energetic, we adopt a “tilted-gravity” formulation. Specifically, we tilt the terrain by a small slope angle $\theta_{\text{slope}}$, so that gravity acquires a constant component along the direction of motion, and we model Coulomb friction at the load–ground interface with coefficient $\mu_{\text{ground}}$. The resulting downslope component of gravity is mathematically equivalent to a constant horizontal pulling force and can be interpreted as a minimal lumped representation of the net directional work exerted by the dog through the tugline; this effective pull compensates frictional and other dissipative losses and thereby enables sustained periodic motion without introducing additional limb-level actuation parameters. All parameters are summarized in Table 1 and are normalized by the total system mass $m$, gravitational acceleration $g$, and resting leg length $l_{0}$. The initial values of these parameters are also reported in Table 1, where the quadruped model parameters were inherited from the model in Ding2024RALSymmetry, which was identified from animal data.
@@
Table 1: Key model parameters used in the quadrupedal SLIP model with load. All quantities are normalized using total mass $m$, gravity $g$, and resting leg length $l_{0}$ GanAllCommonBipedal.
Symbol Description Units Initial Value
$M_{\rm torso}$ Torso mass $m$ 1
$m_{\rm foot}$ Mass of each foot $m$ 0
$J_{\rm torso}$ Torso pitching inertia $ml_{0}^{2}$ 1.1
$k_{\rm leg}$ Leg spring stiffness $mg/l_{0}$ 10
$r_{\rm leg}$ Stiffness ratio between hind/fore legs $\cdot$ 1
$k_{\mathrm{swing},i}^{n}$ Torsional stiffness of swing‐leg spring $mgl_{0}$ 20
$l_{l}$ Resting leg length $l_{0}$ 1
$l_{b,H}$, $l_{b,F}$ Distance from CoM to hip/shoulder $l_{0}$ 0.5
$M_{\rm load}$ Load mass $m$ 0.1
$k_{\rm tug}$ Tugline (unilateral spring) stiffness $mg/l_{0}$ 10
$l_{t}$ Tugline resting length $l_{0}$ 2
$\mu_{\text{ground}}$ Friction coefficient of ground contact $\cdot$ 0.1
$\theta_{\text{slope}}$ Slope angle of running surface $\cdot$ 0.1